Find the slope of the curve at the given point P and an equation of the tangent line at P.


y = -3 - x3, (1, -4)


1
Expert's answer
2021-04-29T16:19:14-0400

Let the slope of the function "y = -3 - x^3" be "m_1"



"m_1 = \frac{dy}{dx}=-3x^2"

at P(1,-4),


"m_1 = -3(1)^2 = -3(1)=-3"

The tangential line to the curve at P(1,-4) is parallel to the curve, thus


"m_1=m_2"

where "m_2" is the slope of the equation of tangential line.


The equation of the tangent line can be obtained thus:


"m_2 = \frac{y-y_1}{x-x_1}"

where:


"x_1=1; y_1=-4"

"m_2 = \frac{y-(-4)}{x-1}=\frac{y+4}{x-1}"

since "m_2=3" , then


"3=\frac{y+4}{x-1}\\
3(x-1)=y+4\\
3x-3=y+4\\
3x-y =4+3\\
3x-y=7 \implies y=3x-7"

is the required tangential equation.